OpenStudy (anonymous):

f(x) = −2x + 4, find x if f(x) = 20. would the answer be -36?

OpenStudy (anonymous):

No so when F(x)=20 you will get 20=-2x+4 then solve for x

OpenStudy (anonymous):

haha 1/4

OpenStudy (anonymous):

zED is right

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

um im confused cuz i did it like you said but i got +8 nd in the choices there is only -8?

OpenStudy (anonymous):

what other answers were given? But it should be -8

OpenStudy (anonymous):

20=-2x+4 16=-2x x=-8

OpenStudy (anonymous):

ohh my bad i got it right i forgot the - sign but there is one i did the same as you said nd it didnt give me the right answer

OpenStudy (anonymous):

f(x) = 2(x + 6), find x if f(x) = 22.

OpenStudy (anonymous):

i got 6

OpenStudy (anonymous):

close double check it, you're a little off

OpenStudy (anonymous):

haha -1

OpenStudy (anonymous):

okay i did 22= 2(x+6) 22=2x +12 -12 -12 _____________ 12 =2x 12/2 =2x/2 6 = x ?

OpenStudy (anonymous):

okay when you minus 12 from 22 you will get 10

OpenStudy (anonymous):

2x=10

OpenStudy (anonymous):

ohh lordyy dumb mistake my bad but thanks for the help

OpenStudy (anonymous):

\[\cup\] when its that symbol again what does that mean?

OpenStudy (anonymous):

it means the union of. E.g. A ∪ B means the set of those elements which are either in A, or in B, or in both.

OpenStudy (anonymous):

so howcan you tell when it is a funtion?

OpenStudy (anonymous):

@shinigamiLm can you do me a hudge favor and grow up just a lil cuz there are people who use this website for help nd others that act like ur 5 so please if ur gonna comment on my questions with unessary comments just hold it it cuz i use this website to help me cuz i need ppl to explain things step by step im confused most of the time cuz math is not my strongest subject so can you please just let me learn nd get the help i need without ur confusion thanks bunches i hate to come off as rude but thats just ignorant well have a merry christmas nd happy new year if you celebrate :)